OpenStudy (anonymous):

is y=2^0.1x growth or decay

OpenStudy (ybarrap):

take the slope: say between x = 1 and x = 2: (y(2) - y(1) / (2 -1) = (2^(0.1*2) - 2^(0.1*1)) / 1 > 0 since 0.1*2 - 0.1.1 = 0.1(2 -1) > 0 Since the slope is positive, what can you say about y itself? If you track your money between year 2 and year 1 and the slope is positive, is the amount of money saved going up if the slope is positive between those years (i.e. (money next year - money this year)/ 1 year > 0)